# Laser mode-hopping assisted all-optical single beam pulsed atomic magnetometer

### v1

@@ IV Experimental results and discussion
Changes in the LD injection current lead to concurrent changes in laser beam intensity and detuning, consequently affecting both the Faraday rotation angle and the spin relaxaton rate. In order to examine these parameters as a function of the LD injection currents, we obtained FID signals by chainging the LD current DC offset at the probe stage while current modulation at the pumping stage is fixed. Because the spin polarization remains unchanged (due to fixed current modulation at the pumping stage), the FID signal reflects the atomic response on the LD current DC offset. Fig. 3 (a) depicts the absorption spectra of 87Rb in the atomic vapor cell. Fig. 3 (b) shows the laser frequency detuning as a function of LD injection currents. A huge shift on the laser frequency occurs around 67.1 mA. The relative size of Faraday rotation angles and transverse spin realxation rates as a function of LD injection currents is shown at Fig. 3 (c). At the current lower than the mode-hop current, the laser becomes near-resonant so that a large Faraday rotation angle and a fast spin relaxation are induced. The LD current higher than the mode-hop current, conversely, shifts the laser frequency into off-resonance, leading to a relatively small Faraday rotation angle and a smaller spin relaxation rate. The maximum difference in spin relaxation rates between on- and off-resonance was approximately 1036 /s, which represents the maximum optical pumping rate. The minimum spin relaxation rate in off-resonance is 2392  /s when the $1/e^{2}$ beam diameter is 10 mm which is the internal size of the atomic vapor cell. When a beam diameter is small, Rb-Rb spin exchange collisions act as the main spin relaxation mechanism. A polarized atomic spin in the boundary between the bright region and the dark region is easily depolarized due to a spin exchange collision with an unpolarized atomic spin from the dark region. With a beam diameter of 1.5 mm, the minimum spin relaxation rate was measured as 4429  /s which is very close to 4304  /s, the 87Rb spin-exchange relaxation rate at the cell temperature of 98 C gibbs1967spin . It means that the optical pumping can be effectively turned off by using laser mode hopping. Fig. 3 (c) shows the optical power of the laser beam after passing through the atomic vapor cell. The absorption is minimized at the off-resonance condition (laser current above the mode-hop current), leading to a higher FID amplitude $V_{0}$ in spite of the slight degradation of the Faraday rotation angle.
@@
We measured the initial amplitude of the FID as a function of the LD current DC offset and the current modulation depth as shown in Figure. 4. Fig. 4 (a) shows the measured relative FID amplitude. We divide Fig. 4 (a) into the four regions depending on whether the current modulation in the pumping stage includes the laser mode-hop current and whether the current in the probing stage is larger(or smaller) than the laser mode-hop current. The black dashed lines in Fig. 4 (a)-(c) indicate the boundary where the current modulation in the pumping stage contains the laser mode-hop current. The black solid line depicts the mode-hop current, 67.1 mA. The laser frequeny in the probing stage is off-resonant (or on-resonant) in the region to the right (or left) of the black solid line. We note that the experimental data shown in Fig. 4 (a) reflects the combined effect of the paramagnetic Faraday rotation, beam absorption and the discontinuous optical pumping.
@@
In the region 1, current modulation in the pumping stage excludes the laser mode-hop current so that the continuous optical pumping profile with a small $R_{1}$ causes the small $S_{0}$. Additionally, the current in the probing stage brings the laser frequency into on-resonance, resulting in a small beam intensity due the strong absorption. These lead to the small FID amplide as depicted by the black line in Fig. 4 (d). In the region 2, current modulation in the pumping stage does not contain the mode-hop current like the region 1. However, $R_{1}$ of the region 2 is smaller than region 1 as shown in the insets of Fig. 4 (d). Hence, $S_{0}$ of the region 2 is expected to be smaller than that of the region 1. On the other hand, the laser frequency at the probing stage is off-resonant so that the spin relaxation rate is much smaller than that of region 1. The resulting FID signal is depicted by the green line in Fig. 4 (d), whose amplitude is slightly smaller than region 1 (black line). In the region 3, the mode-hop current is included in the current modulation, resulting in a highly discontinuous modulation profile. A large $S_{0}$ is produced and a large Faraday rotation occurs at the probing stage. In this region laser frequency at the probing stage is on-resonant. The absorption of the laser beam, therefore, results in a degradation of FID amplitude, as shown in red line in Fig. 4 (d). The maximum FID amplitude is shown in the region 4, as depicted by the blue line in Fig. 4 (d). A large $S_{0}$ is induced because the current modulation including the laser mode-hop current realize a discontinous optical pumping profile with a large $R_{1}$. The off-resonant beam minimizes absorption at the probing stage so that a large FID amplitude is induced compared to region 3. Moreover, the spin relaxation rate of the region 4 is slower compared to that of region 3, which is beneficial in terms of magnetometer sensitivities.
@@
To compare the experimental data with the theory, Eq. (2) is calculated using $R_{0}$ and $R_{1}$ extracted from the $R(t)$ profile shown at the top panel of Fig. 4 (b) and (c). The $R(t)$ profile in the top panel of Fig. 4 (b) is obtained by subtracting the minimum relaxation rate from the measured spin relaxation rate shown at Fig. 3 (c). Fig. 4 (b) shows the calculated relative FID amplitude using the triangular function like $R(t)$ profile in the top panel of Fig. 4 (b). Fig. 4 (c) is derived from the fictious step-function like $R(t)$ profile given in the top panel of Fig. 4 (c). Both the experimental result and the semi-theoretical results seem to have similar tendency. But the details are slightly different from each other. Figure 5 shows maximum FID amplitudes as a function of modulation depth and the LD current DC offsets with the maximum FID amplitude in the experiment (Fig. 4 (a)) and two semi-theoretical calculations (Fig. 4 (b) and (c) ). Because two semi-theoretical results and the experimental result use the same Faraday rotation profile and the same laser beam absorption profile, the differences between them are mainly attributed by discontinuous optical pumping.
@@
Figure 5: (a) Maximum FID amplitude as a function of modulation depth. The black dots represent the experimental data, derived from Fig. 4 (a). The red line and the blue line depict the semi-theoretical results extracted from Fig. 4 (b) and Fig. 4 (c), respectively. The experimental data(black dots) is more robust to the change of modulation depth compared to the semi-theoretical result(red line) where the optical pumping profile is a triangular function. The semi-theoretical result (blue line) where the optical pumping profile is a step function shows that the maximum ampltiude of FID increases with increasing modulation depth. (b) LD current DC offset with the maximum FID amplitude as a function of modulation depth.
@@
Comparing experimental data (black dots) and semi-theoretical result with the measured triangular function like optical pumping profile (red line) in Fig. 5 (a), the maximum FID amplitude decreases moderately as the modulation depth grows while $S_{0}$ derived by Eq. (2) fastly decreases. It implies that the measured triangular function like optical pumping profile used in the top panel of Fig. 4 (b) does not fully describe the mode-hopping asssited optical pumping. This discrepancy between the semi-theoretical result and the experimental result needs to be studied in the future. Here, we describe its potential origins. First, we note that the optical pumping rate profile shown in Fig. 4 (b) is calculated from the transverse spin relaxation rates extracted from the FID signals as a function of the LD current DC offset in the probing stage. The FID signals are affected by the constant laser frequency. However, because the laser frequency sweeps in the pumping stage, the optical pumping rate in the pumping stage is slightly different from the optical pumping rate derived from the FID signals as a function of the LD current DC offset in the probing stage. In particular, when the ground hyperfine structure is considered, the optical pumping rates of the $F_{g}=1$ state and the $F_{g}=2$ change with times because of the laser frequency sweeping. The transverse spin along the $\hat{x}$-direction can be represented by $\langle S_{x}\rangle=(\langle F_{x}^{I+1/2}\rangle-\langle F_{x}^{I-1/2}\rangle)/(2I+1)$ where $I$ represents the nulcear spin, $I=3/2$ for 87Rb. To maximize spin polarization $S_{0}$ of 87Rb, the $F_{g}=2,~{}m_{F_{g}}=2$ state (or $m_{F_{g}}=-2$) should be mainly populated. For this, the hyperfine repumping of the $F_{g}=1$ state is required li2022repumping ; schultze2015improving . Due to the large pressure broadening of a order of GHz, the laser at the pumping stage pumps both the $F_{g}=1$ and $F_{g}=2$ states with different optical pumping rates. The different time-varying optical pumping rates for the $F_{g}=1$ and the $F_{g}=2$ states have to be taken into account to describe the mode-hopping assisted optical pumping more accurately. Second, the spatial inhomogeneity of the optical pumping profile needs to be considered. Due to light absorption, the optical pumping beam attenuates along the beam propagation direction so that the optical pumping profile is spatially inhomogeneous. The measured optical pumping profile shown in the top panel of Fig. 4 (b) is a kind of spatially averaged values. The analysis considering the spatially inhomogeneous optical pumping profile will give more accurate results. Additionally, we obtain the semi-theoretical result with the step function like optical pumping profile as depicted by the blue solid line in Fig. 5 (a). In contrast to the experimental result and the semi-theoretical result with the triangular function like optical pumping profile, the relative FID amplitude increases with increasing modulation depth. Interestingly, the experimental data is between two semi-theoretical results. We can conjecture that the real optical pumping profile is between the triangular function and the step function.
We estimate the sensitvity of our magnetometer based on the laser mode-hopping assisted optical pumping. We measured Larmor frequencies at a repetiton rate of 1 kHz by using the commercial frequency counter. The measured Larmor frequencies were converted into the magnitude of the magnetic field by diving the Larmor frequencies by the gyromagnetic ratio of 87Rb. Finally, the amplitude spectral density (ASD) of the magnitude of the magnetic field was obtained. The inset of Figure 6 (a) shows averaged ASD of 1000 shots of the FID signal when the beam intensity after passing through the atomic vapor cell was 0.9 mW. Fig. 6 (a) shows the sensitivity of our magnetometer obtained by measuring almost 5 sec. We acheived a sensitivity of 5.57 pT/Hz1/2, mainly limited by the performance of the frequency counter. We note that a sensitivity of 3.77 pT/Hz1/2 was obtained in the laser power of 1.4 mW where the FID signal was partially a square-like wave due to the saturation of the BPD.
The measured sensitivity is mainly limited by the amplitude noise in the frequency counter. Considering a damped oscillation with a RMS amplitude noise $\delta y$, the amplitude noise is converted into a frequency error via the slew rate at the threshold. When the frequency is estimated by averaging the time intervals between the zero-crossing time stamps (mean frequency model of frequency counter), the deviation of the magnetic field error $\rho_{fc}$ averaged over the measurement time of $T_{m}$ is written by
$\rho_{fc}=\frac{\delta y}{\gamma V_{0}\sqrt{f_{BW}}T_{m}}\sqrt{1+e^{2\Gamma_{re}T_{m}}}.$ (5)
Here, $V_{0}$ is the FID ampliutde and $f_{BW}$ is the bandwidth of the magnetometer. The total rms amplitude noise $\delta y$ is given by $\delta y=\sqrt{\delta^{2}y_{fc}+\delta^{2}y_{ph}}$ where $\delta y_{fc}$ is the amplitude noise of the frequency counter, 3500 $\mu$$\textrm{V}_{\textrm{rms}}$ and $\delta y_{ph}$ is the RMS photon shot-noise. The calculated magnetic field noise is shown at Fig. 6 (b) as the black dashed line which shows good agreement to the measured magnetometer noise floor.
@@
Our frequency measurement using the commercial frequency counter is not optimal. The potentially achievable sensitivity can be estimated by the Cramer-Rao lower bound (CRLB) calculated by a shot of FID signal gemmel2010ultra . The Cramer-Rao lower bound (CRLB) is the sensitivity reference for an atomic magnetometer based on measurement of the Larmor frequency of the decaying FID signal. The CRLB is given by
@@
where $\rho_{\omega}$ is the ASD level of the BPD signal at the Larmor freuqency. The coefficient $C_{\omega}$ which depends on the transveral spin relaxation rate $\Gamma_{re}$ is given by
Figure 6: (a) Noise spectral density of our magnetometer. The black line is the magnetometer noise floor of 5.57 pT/Hz1/2 which is averaged over 10 Hz to 500 Hz when the commercial frequency counter is applied. The red dashed line shows a CRLB of 191 fT/Hz1/2 which is calculated from Eq. (6). The inset shows the ASD of the FID signal which is used for the calculation of the CRLB. The red line depicts the ASD when $\Delta w=|w-w_{0}|\simeq 0$. The black line depicts the ASD when $\Delta w=100$ kHz, which determines $\rho_{\omega}$. A small 85Rb peak is shown at 143 kHz. The resolution bandwidth of the ASD is 3.9 kHz which is limited by the magnetometer bandwidth. The low frequency ASD includes an internal electronic noise arising from the spectrum analyzer(EXA Signal Analyzer N9010A). The photon shot-noise level is depicted by the dashed line for comparison. The beam power after passing through the atomic vapor cell was 0.9 mW. (b) Magnetometer noise floor and CRLB as a function of beam power after passing thorugh the atomic vapor cell. The black dots depict the measured magnetometer noise floor using the commerical frequency counter. The black dashed line shows the calculated noise level from mean frequency model (Eq. (5)). The red dots show the CRLB calculation based on the ASD measurement, which represent the reachable sensitivity of the magnetometer (Eq. (6)). The red dashed line shows the photon shot-noise limited CRLB calculation.
@@
where $r=1/(\Gamma_{re}T_{m})$. The CRLB calculated by Eq. (6) represents the non magnetic-oriented noise induced sensitivity. When the beam intensity after absorption is 0.9 mW, $\Gamma_{re}=2358$ /s, $T_{m}=460$ $\mu$s, $\rho_{\omega}=2.6$ $\mu$V/Hz1/2, $V_{0}=8.5$ V, and $f_{BW}=500$ Hz, we obtain the CRLB of 191 fT/Hz1/2. Fig. 6 (b) shows the measured magnetometer noise level and CRLB for various laser power after absorption throughout the vapor cell. The red points show the CRLB calculation assuming that $\rho_{\omega}$ is limited to the photon shot-noise which is given by $\sqrt{eG^{2}RI_{0}}$. Here, $G$ is the transimpedence gain, $R$ is the reponsivity of the BPD and $e$ is the classical electron radius. The CRLB calculation shows that our magnetometer is potentially capable of reaching a sensitivity of a sub pT/Hz1/2 by revising the frequency counting method.
We note that the frequency shift of 20 GHz and the laser frequency in the pumping stage which is moderately detuned from the resonance are not optimal. In order to mainly populate the $F_{g}=2,~{}m_{F_{g}}=2$ state for maximizing $S_{0}$, the on-resonance laser frequency in the pumping stage should be close to the $F_{g}=1,2$ resonances. Also the laser freuqency needs to be far-detuned from the resonance periodically for synchronous optical pumping to obtain a large optical pumping modulation depth. A careful laser cavity adjustment in laser diode fabrication, leading to an optimal frequency shift of mode-hopping and the on-resonance laser frequency very close to the resonance, will be helpful to increase the sensitivity of magnetometer. In addition, multi-pass probing will increase the sensitivity because $V_{0}$ is proportional to the interaction length li2011optical .

### v2

@@ IV Experimental results and discussion
Changes in the LD injection current lead to concurrent changes in laser beam intensity and detuning, consequently affecting both the Faraday rotation angle and the spin relaxation rate. To examine these parameters as a function of the LD injection currents, we obtained FID signals by changing the LD current DC offset at the probe stage, while current modulation at the pumping stage remained fixed. Because the spin polarization is unchanged (due to fixed current modulation at the pumping stage), the FID signal reflects the atomic response on the LD current DC offset. Fig. 3 (a) depicts the absorption spectra of 87Rb in the atomic vapor cell. Fig. 3 (b) shows the laser frequency detuning as a function of LD injection currents. A huge shift in the laser frequency occurs at around 67.1 mA.
@@
The relative size of Faraday rotation angles and transverse spin relaxation rates as a function of LD injection currents is shown in Fig. 3 (c). At a current lower than the mode-hop current, the laser becomes nearly resonant so that a large Faraday rotation angle and a fast spin relaxation are induced. Conversely, with an LD current higher than the mode-hop current, the laser frequency shifts into off-resonance, leading to a relatively small Faraday rotation angle and a smaller spin relaxation rate. The maximum difference in spin relaxation rates between on- and off-resonance was approximately 1036 /s, which represents the maximum optical pumping rate. The minimum spin relaxation rate in off-resonance is 2392  /s when the $1/e^{2}$ beam diameter is 10 mm, which is the internal size of the atomic vapor cell. When the beam diameter is small, Rb-Rb spin exchange collisions act as the main spin relaxation mechanism. A polarized atomic spin in the boundary between the bright region and the dark region is easily depolarized by a spin exchange collision with an unpolarized atomic spin from the dark region. With a beam diameter of 1.5 mm, the minimum spin relaxation rate was measured as 4429  /s which is very close to 4304  /s, the 87Rb spin-exchange relaxation rate at the cell temperature of 98 C gibbs1967spin . This means that the optical pumping can be effectively turned off by using laser mode hopping. Fig. 3 (c) shows the optical power of the laser beam after passing through the atomic vapor cell. The absorption is minimized in the off-resonance condition (laser current above the mode-hop current), leading to a higher FID amplitude $V_{0}$ despite the slight degradation in the Faraday rotation angle.
@@
We measured the initial amplitude of the FID as a function of the LD current DC offset and the current modulation depth as shown in Figure. 4. Fig. 4 (a) shows the measured relative FID amplitude. We divide Fig. 4 (a) into the four regions, depending on whether the current modulation in the pumping stage includes the laser mode-hop current and whether the current in the probing stage is larger (or smaller) than the laser mode-hop current. The black dashed lines in Fig. 4 (a)-(c) indicate the boundary where the current modulation in the pumping stage contains the laser mode-hop current. The black solid line depicts the mode-hop current, 67.1 mA. The laser frequency in the probing stage is off-resonant (or on-resonant) in the region to the right (or left) of the black solid line. We note that the experimental data shown in Fig. 4 (a) reflects the combined effect of the paramagnetic Faraday rotation, beam absorption and the discontinuous optical pumping.
@@
In region 1, the current modulation in the pumping stage excludes the laser mode-hop current so that the continuous optical pumping profile with a small $R_{1}$ causes a small $S_{0}$. Additionally, the current in the probing stage brings the laser frequency into on-resonance, resulting in a small beam intensity due the strong absorption. These factors lead to the small FID amplitude as depicted by the black line in Fig. 4 (d). In region 2, the current modulation in the pumping stage does not contain the mode-hop current, as in region 1. However, the value of $R_{1}$ in region 2 is smaller than that in region 1 as shown in the insets of Fig. 4 (d). Hence, $S_{0}$ of the region 2 is expected to be smaller than that of region 1. On the other hand, the laser frequency in the probing stage is off-resonant, resulting in a spin relaxation rate that is much smaller than that in region 1. The resulting FID signal is depicted by the green line in Fig. 4 (d), whose amplitude is slightly smaller than region 1 (black line). In region 3, the mode-hop current is included in the current modulation, resulting in a highly discontinuous modulation profile. A large $S_{0}$ is produced and a large Faraday rotation occurs in the probing stage. In this region the laser frequency in the probing stage is on-resonance. The absorption of the laser beam, therefore, results in a degradation in FID amplitude, as shown by the red line in Fig. 4 (d). The maximum FID amplitude is in region 4, as shown by the blue line in Fig. 4 (d). A large $S_{0}$ is induced because the current modulation, including the laser mode-hop current, results in a discontinuous optical pumping profile with a large $R_{1}$. The off-resonant beam minimizes absorption in the probing stage, inducing a large FID amplitude compared to region 3. In addition, the spin relaxation rate in region 4 is slower than in region 3, which is beneficial for magnetometer sensitivity.
@@
To compare the experimental data with the theory, Eq. (2) is calculated using $R_{0}$ and $R_{1}$ extracted from the $R(t)$ profile shown at the top panel of Fig. 4 (b) and (c). The $R(t)$ profile in the top panel of Fig. 4 (b) is obtained by subtracting the minimum relaxation rate from the measured spin relaxation rate shown in Fig. 3 (c). Fig. 4 (b) shows the calculated relative FID amplitude using the triangular function like $R(t)$ profile in the top panel of Fig. 4 (b). Fig. 4 (c) is derived from the fictitious step-function like $R(t)$ profile given in the top panel of Fig. 4 (c). Both the experimental result and the semi-theoretical results seem to have similar tendencies. But the details are slightly different from each other. Figure 5 shows the maximum FID amplitudes as a function of the modulation depth and the LD current DC offsets with the maximum FID amplitude in the experiment (Fig. 4 (a)) and two semi-theoretical calculations (Fig. 4 (b) and (c) ). Since the two semi-theoretical results and the experimental result use the same Faraday rotation profile and the same laser beam absorption profile, the differences between them are mainly attributed to the discontinuous optical pumping.
@@
Figure 5: (a) Maximum FID amplitude as a function of modulation depth. The black dots represent the experimental data, derived from Fig. 4 (a). The red line and the blue line depict the semi-theoretical results extracted from Fig. 4 (b) and Fig. 4 (c), respectively. The experimental data (black dots) is more robust to the change in modulation depth compared to the semi-theoretical result (red line) where the optical pumping profile is a triangular function. The semi-theoretical result (blue line) where the optical pumping profile is a step function shows that the maximum amplitude of FID increases with increasing modulation depth. (b) LD current DC offset with the maximum FID amplitude as a function of modulation depth.
Comparing the experimental data (black dots) and semi-theoretical result with the measured triangular function like optical pumping profile (red line) in Fig. 5 (a), the maximum FID amplitude decreases moderately as the modulation depth grows while $S_{0}$ derived by Eq. (2) rapidly decreases. This implies that the measured triangular function like optical pumping profile used in the top panel of Fig. 4 (b) does not fully describe the mode-hopping assisted optical pumping.
@@
This discrepancy between the semi-theoretical result and the experimental result needs to be studied in the future. Here, we describe its potential origins. First, we note that the optical pumping rate profile shown in Fig. 4 (b) is calculated from the transverse spin relaxation rates extracted from the FID signals as a function of the LD current DC offset in the probing stage. The FID signals are affected by the constant laser frequency. However, because the laser frequency sweeps in the pumping stage, the optical pumping rate in the pumping stage is slightly different from the optical pumping rate derived from the FID signals as a function of the LD current DC offset in the probing stage. In particular, when the ground hyperfine structure is considered, the optical pumping rates of the $F_{g}=1$ state and the $F_{g}=2$ change with times because of the laser frequency sweeping. The transverse spin along the $\hat{x}$-direction can be represented by $\langle S_{x}\rangle=(\langle F_{x}^{I+1/2}\rangle-\langle F_{x}^{I-1/2}\rangle)/(2I+1)$ where $I$ represents the nuclear spin, $I=3/2$ for 87Rb. To maximize spin polarization $S_{0}$ of 87Rb, the $F_{g}=2,~{}m_{F_{g}}=2$ state (or $m_{F_{g}}=-2$) should be mainly populated. For this, the hyperfine repumping of the $F_{g}=1$ state is required li2022repumping ; schultze2015improving . Due to the large pressure broadening on the order of GHz, the laser at the pumping stage pumps both the $F_{g}=1$ and $F_{g}=2$ states with different optical pumping rates. The different time-varying optical pumping rates for the $F_{g}=1$ and the $F_{g}=2$ states have to be taken into account to describe the mode-hopping assisted optical pumping more accurately.
Second, the spatial inhomogeneity of the optical pumping profile needs to be considered. Due to light absorption, the optical pumping beam is attenuated along the beam propagation direction so that the optical pumping profile is spatially inhomogeneous. The measured optical pumping profile shown in the top panel of Fig. 4 (b) is a kind of spatially averaged values. The analysis considering the spatially inhomogeneous optical pumping profile will give more accurate results. Additionally, we obtained the semi-theoretical result with the step-function-like optical pumping profile, as depicted by the blue solid line in Fig. 5 (a). In contrast to the experimental result and the semi-theoretical result with the triangular function like optical pumping profile, the relative FID amplitude increases with increasing modulation depth.
Interestingly, the experimental data lie between two semi-theoretical results. We can conjecture that the real optical pumping profile lies between the triangular function and the step function.
We estimate the sensitivity of our magnetometer based on the laser mode-hopping assisted optical pumping. In order to further decrease the effect of the unwanted external magnetic field and to increase the uniformity of the magnetic field, we applied the 4 layered $\mu$-metal shield (MS-1, Twinleaf) instead of the small home-made 2 layered $\mu$-metal shield which were used to test the functionality of the laser mode-hopping. After applying the bias magnetic field along the z-axis of the magnetic shield, we measured Larmor frequencies at a repetition rate of 1 kHz by using the commercial frequency counter. The measured Larmor frequencies were converted into the magnitude of the magnetic field by diving the Larmor frequencies by the gyromagnetic ratio of 87Rb. Finally, the amplitude spectral density (ASD) of the magnitude of the magnetic field was calculated as shown in Fig. 6 (a). The magnetic field noise floor level is 0.6 pT/Hz1/2 at a magnetic field of 27 $\mu$T, as depicted by the red dashed line in Fig. 6 (a).
In order to observe the magnetic field noise floor limited by the frequency counting method based on the linear regression of zero-crossing times, we obtained the ASDs as a function of Larmor frequencies. The Larmor frequency changes depending on the current flowing in the bias coil. When the frequency is extracted from the linear regression of zero-crossing times, the magnetic field noise can be calculated as
$\rho_{fc}=\frac{12\delta y~{}f_{0}}{2\pi\gamma V_{0}N^{3}\sqrt{f_{BW}}}\sqrt{\sum_{j}^{N}{(j-\frac{\sum j}{N})^{2}}e^{2j\Gamma_{2}/f_{0}}},$ (5)
where $j$s are integer numbers and $N$ is the total number of the zero-crossings, given by $N=f_{0}T_{m}$ where $f_{0}$ is the Larmor frequency and $T_{m}$ is the measurement time. The total rms amplitude noise $\delta y$ is given by $\delta y=\sqrt{\delta^{2}y_{fc}+\delta^{2}y_{ph}}$ where $\delta y_{fc}$ is the amplitude noise of the frequency counter, 350 $\mu$$\textrm{V}_{\textrm{rms}}$ and $\delta y_{ph}$ is the RMS photon-shot-noise. The amplitude spectral density of the photon-shot-noise is given by $\sqrt{2eG^{2}RP_{0}}$, where $G$ is the transimpedence gain ($G=10^{5}$ V/A for a high-Z load), $R$ is the responsivity of the BPD, $P_{0}$ is the optical power in front of the polarimeter, and $e$ is the electron charge. When the total power of the probe beam after passing through the atomic gas cell was 4 mW, the measured ASD including the probe electronic noise and the photon-shot-noise was 2.0 $\mu$V/Hz1/2. It gave a RMS voltage of 1400 $\mu$$\textrm{V}_{\textrm{rms}}$ in the fixed noise band width of $N_{BW}=500$ kHz. The magnetic field noise originating from the electronic noise and the photon-shot-noise, as a result of Eq. (5), is represented by the blue dashed line in Fig. 6 (b). In opposition to the result of Eq. (5),the measured noise floor increases with the averaged current.
In order to investigate such linear dependence of the magnetic field noise on the averaged current, we considered the magnetic field noise driven by the current noise of the current supply (CSUA-300, Twinleaf). We measured the current noise of the current supply by using a signal analyzer (N9010A, Keysight). The current noise was linearly proportional to the averaged current $I\right>$ and it can be represented as $\delta I=\alpha+\beta\left<I\right>$ where $\alpha=0.6$ (0.08) nA/Hz1/2 and $\beta=11.3$ (0.8) nA/Hz1/2/A. By considering the coil calibration constant, 130 $\mu$T/A, the magnetic field noise due to the current noise of the current supply can be estimated by $\delta B_{cc}=$ 1.47 pT/Hz1/2/A ×$I\right>$ + 0.078 pT/Hz1/2 . The magenta dashed line in Fig. 6 (b) depicts the magnetic field noise due to the current noise from the current supply. The squared sum combination of the photon-shot-noise and the magnetic field noise from the current noise of the current supply, $\delta B=\sqrt{\rho^{2}_{fc}+\delta B_{cc}^{2}}$, is represented by the black dashed line in Fig. 6 (b). It shows a good agreement with the experimentally measured magnetic field noise floor.
@@
Our frequency measurement using the commercial frequency counter was not optimal. The potentially achievable sensitivity can be estimated by the CRLB, calculated using a shot of the FID signal gemmel2010ultra . The CRLB is the sensitivity reference for an atomic magnetometer based on measurement of the Larmor frequency of the decaying FID signal. The CRLB is given by
@@
where $\rho_{\omega}$ is the ASD level of the BPD signal at the Larmor frequency. The coefficient $C_{\omega}$ which depends on the transversal spin relaxation rate $\Gamma_{re}$ which is given by
Figure 6: (a) Amplitude spectral density of the magnetic field noise at a bias field of 27 $\mu$T. The red dashed line represents a magnetic noise floor of 0.6 pT/Hz1/2. (b) Noise as a function of the Larmor frequencies. The blue dashed line depicts the calculation for frequency counting using Eq. (5). The magenta dashed line represents the magnetic field noise generating from the bias magnetic field coil driven by the current supply with a current noise of 11.3 nA/Hz1/2/A. The green dashed-dot lines shows the Cramér-Rao lower bound (Eq. (6)), 0.131 pT/Hz1/2 in our experimental conditions.
@@
where $r=1/(\Gamma_{re}T_{m})$. The CRLB calculated by Eq. (6) represents the non-magnetic oriented noise induced sensitivity. After optimization, the parameters for calculating the CRLB are given as $\Gamma_{re}=1670$ /s, $T_{m}=650$ $\mu$s, $\rho_{\omega}=2$ $\mu$V/Hz1/2, $V_{0}=8$ V, and $f_{BW}=500$ Hz in our experiments using the four layered $\mu$ shield. We obtain the CRLB of 131 fT/Hz1/2. The CRLB calculation shows that our magnetometer is potentially capable of reaching a sensitivity of a sub pT/Hz1/2 by revising the frequency counting method; the noise bandwidth is adjusted to the Larmor frequency, i.e., $N_{BW}\simeq f_{0}$.
We note that the frequency shift of 20 GHz, and the laser frequency in the pumping stage which is moderately detuned from the resonance, are not optimal. In order to mainly populate the $F_{g}=2,~{}m_{F_{g}}=2$ state to maximize $S_{0}$, the on-resonance laser frequency in the pumping stage should be close to the $F_{g}=1,2$ resonances. Also the laser frequency needs to be periodically far-detuned from the resonance for synchronous optical pumping, to obtain a large optical pumping modulation depth. A careful laser cavity adjustment in the laser diode fabrication, leading to an optimal frequency shift of mode-hopping and the on-resonance laser frequency very close to the resonance, will be helpful to increase the sensitivity of the magnetometer. In addition, multi-pass probing will further increase the sensitivity because $V_{0}$ is proportional to the interaction length li2011optical .